Question1 1 pts Find the marginal utility of an additional unit of good a for a consumer with utility function u (x, y) yi when their current consumption level i 2 and y 16. HTML Editor
Answer #1

Marginal utility (MU) of an additional unit of good x is d(u(x,y)) / dx

u(x,y) = x2/3 y1/3

MU of good x (MUx) = (2/3)* x-1/3 y1/3

= (2/3) *(y/x)1/3

given, x = 2 and y = 16

MUx = (2/3) (16/2)1/3

MUx = 4/3 = 1.33